Mike has dimes and quarters. He has 18 coins. Mike has a total of $4.20. How many of each does he have?
Answer by ewatrrr(24785)

x the number of Quarters
CENTS makes sense
25x+ 10(18-x) = 420
x = 240/15
x = 16, number of Quarters. 2 Dimes 18-16
And...checking
25*16 + 2*10 = 400 + 20 = 420